8 September

Participants

name  place  e-mail
Bernd Ackermann(D)  Stuttgart  ackermann@mathematik.uni-stuttgart.de
Jan Adrienssen (B)  Antwerp  adri@ruca.ua.ac.be
Anne-Marie Aubert(F)  Paris  Anne-Marie.Aubert@ens.fr
Karin Baur (Ch)  Basel  baur@math.unibas.ch
Jens Bender (D)  Wuppertal  bender@uni-wuppertal.de
Rafael Bocklandt (B)  Antwerp  rbockl@hwins.uia.ac.be
G. Bobinski (P)  Torun  gregbob@mat.uni.torun.pl
Giovanna Carnovale (I)  Antwerp  gcarno@uia.ua.ac.be
Corina Calinescu (R)  Bucharest  Corina.Calinescu@imar.ro
Christelle Chesne (F)  Duesseldorf  chesne@uni-duesseldorf.de
Anton Cox (UK)  London  A.G.Cox@city.ac.uk
Steve Doty (US)  Chicago  doty@math.luc.edu
Matt Fayers (UK)  Cambridge  mf205@cam.ac.uk
Eddy Goddelle (F)  Amiens  eg-math@u-picardie.fr
Lutz Hille (D)  Hamburg  hille@math.uni-hamburg.de
A. Hubery (UK)  Leeds  ahubery@amsta.leeds.ac.uk
Bernt T. Jensen (N)  Trondheim  berntj@math.ntnu.no
Alastair King (UK)  Bath  masadk@maths.bath.ac.uk
O. Khomenko (Ukr)  Kaiserslautern khomenko@mathematik.uni-kl.de
Dirk Kussin (D)  Paderborn  dirk@mathematik.uni-paderborn.de
Andrea Maffei (I)  Rome  amaffei@mat.uniroma1.it
Dag Madsen(N)  Trondheim  dagma@math.ntnu.no
Robert Marsh (UK)  Leicester  rjm25@mcs.le.ac.uk
Stephan Mohrdieck (D)  Basel  mohrdiec@x4u2.desy.de
H. Nagase(J)  Bielefeld  hnagase@mathematik.uni-bielefeld.de
Adriana Nenciu (R)  Bucharest  adriana.nenciu@imar.ro
Florin Panaite (R)  Bucharest  fpanaite@stoilow.imar.ro
Mike Prest (UK)  Manchester  mprest@fs2.ma.man.ac.uk
Markus Reineke(D)  Wuppertal  reineke@math.uni-wuppertal.de
Konstanze Rietsch (Au)  Cambridge  k.rietsch@dpmms.cam.ac.uk
Raphael Rouquier (F)  Paris  rouquier@math.jussieu.fr
D. Rumynin  Warwick  rumynin@maths.warwick.ac.uk
Olivier Schiffmann (F)  Paris  schiffma@clipper.ens.fr
Jan Schroer(D)  Leeds  jschroer@amsta.leeds.ac.uk
B. Sevenhant (B)  Hasselt  bsevenha@luc.ac.be
Rupinder Kaur Sethi (N)  Oslo  rsethi@math.uio.no
Desmond Sheiham (UK)  Edinburgh  des@maths.ed.ac.uk
James Shepherd (UK)  Bristol  james.shepherd@bris.ac.uk
M. D. Staic (R)  Bucharest  mstaic@pompeiu.imar.ro
Catharina Stroppel (D)  Freiburg  stroppel@sun2.mathematik.uni-freiburg.de
Fabio Stumbo (I)  Ferrara  stumbo@dm.unife.it
L. Csaba Szanto (R)  Cluj  szanto@math.ubbcluj.ro
A. Tudorache (R)  Bucharest  ana tudorache@hotmail.com
Jon Eivind Vatne (N)  Bergen  JonEivind.Vatne@mi.uib.no
Geert van der Weyer (B)  Antwerp  geert.vandeweyer@ufsia.ac.be
Jie Xiao (Cn)  Beijing  jxiao@math.tsinghua.edu.cn
Alexander Zimmermann (D) Amiens  alex@math.jussieu.fr

Invited Speakers H. Derksen, Boston. hderksen@math.mit.edu, V. Ginzburg , Chicago. ginzburg@math.uchicago.edu , L. le Bruyn, Antwerp. lebruyn@hwins.uia.ac.be , G. Lusztig, Boston. gyuri@math.mit.edu , H. Nakajima, Kyoto. nakajima@kusm.kyoto-u.ac.jp , A. Schofield, Bristol. Aidan.Schofield@bristol.ac.uk, G. Zwara, Torun. gzwara@mat.uni.torun.pl

Organizers K. Erdmann, Oxford. erdmann@maths.ox.ac.uk , W. Crawley-Boevey, Leeds. w.crawley-boevey@leeds.ac.uk , C. Geiß, UNAM. christof@math.unam.mx